5. A thin layer of hydrocarbon oil (n0 = 1.40) is formed on the wet road (nw = 1.33). Colourful reflections can be seen from the film when illuminated under white light.

(a) A part of the film thickness is 300 nm. Determine the visible wavelength that can be seen on the part.

(b) If the same thickness of kerosene in (a) is spread on a glass (n = 1.5), determine the visible wavelength that can be seen on the part.


1
Expert's answer
2021-07-12T12:14:15-0400

a)

"\lambda_1=\frac{n_o}{n_w}\lambda=\frac{1.40}{1.33}\cdot 300=315~nm,"

b)

"\lambda_2=\frac{n}{n_w}\lambda=\frac{1.5}{1.33}\cdot 300=338~nm."